### v2

# Weingarten calculus with virtual isometries

Abstract.
In this paper, we develop a novel approach to the Weingarten calculus employing the notion of virtual isometries. Traditionally, Weingarten calculus provides explicit formulas for integrating polynomial functions over compact matrix groups with respect to the Haar measure, yet faces limitations when evaluating high-degree integrals due to the non-invertibility of the associated matrices. We revisit these classical computations from a new perspective: by constructing Haar-distributed matrices via sequences of complex reflections, we derive new recursive structures for the Weingarten functions across different dimensions. This framework leads to two main results: (1) an explicit Weingarten calculus for complex reflections, yielding systematic moment computations for associated rank-one matrices, and (2) a novel convolution formula that connects Weingarten functions in dimension $n$ to those in dimension n1n-1 through the introduction of ascension functions in the symmetric group algebra. Our approach not only provides a unified treatment for unitary groups, but also sheds light on the algebraic and probabilistic aspects of high-degree integral computations. Several examples and applications are presented.
Key words and phrases:
unitary group, Haar measure, Weingarten calculus, virtual isometry, complex reflection
2020 Mathematics Subject Classification:
60B20, 15B52, 28C10
1. Introduction
1.1. Weingarten calculus
Weingarten calculus is a method that allows one to explicitly calculate integrals of polynomial functions on compact groups with respect to the Haar probability measure.
An important case – on which we focus in this paper – is the unitary group $\mathrm{U}(n)$, which consists of all $n\times n$ unitary matrices. This case is important because every compact matrix group is conjugate to a subgroup of the unitary group. Let $U=(u_{ij})_{1\leq i,j\leq n}$ be a Haar-distributed matrix from $\mathrm{U}(n)$. From the properties of the Haar measure, when $k$ and $l$ are different nonnegative integers, the expectation of the following form is zero:
$\mathbb{E}\left[u_{i_{1}j_{1}}u_{i_{2}j_{2}}\cdots u_{i_{k}j_{k}}\overline{u_{i_{1}^{\prime}j_{1}^{\prime}}u_{i_{2}^{\prime}j_{2}^{\prime}}\cdots u_{i_{l}^{\prime}j_{l}^{\prime}}}\right]$
for $i_{p},j_{p},i_{p}^{\prime},j_{p}^{\prime}\in\{1,2,\dots,n\}$ and $p\geq 1$. According to the Weingarten formula [Col03], when $k=l$, this moment can be expressed in the form of a sum over the symmetric group $S_{k}$:
(1.1) $\mathbb{E}\left[u_{i_{1}j_{1}}u_{i_{2}j_{2}}\cdots u_{i_{k}j_{k}}\overline{u_{i_{1}^{\prime}j_{1}^{\prime}}u_{i_{2}^{\prime}j_{2}^{\prime}}\cdots u_{i_{k}^{\prime}j_{k}^{\prime}}}\right]\\ =\sum_{\sigma\in S_{k}}\sum_{\tau\in S_{k}}\delta(i_{\sigma(1)},i_{1}^{\prime})\cdots\delta(i_{\sigma(k)},i_{k}^{\prime})\delta(j_{\tau(1)},j_{1}^{\prime})\cdots\delta(j_{\tau(k)},j_{k}^{\prime})\operatorname{Wg}_{k,n}(\sigma\tau^{-1}).$
Here $\delta(p,q)$ is Kronecker’s delta, and $\operatorname{Wg}_{k,n}$ is a class function on $S_{k}$ known as the (unitary) Weingarten function. See also [Sam80]. The Weingarten function has various representations, but when $k\leq n$, we can immediately obtain the representation
$\operatorname{Wg}_{k,n}(\sigma)=\mathbb{E}\left[u_{11}u_{22}\cdots u_{kk}\overline{u_{\sigma(1)1}u_{\sigma(2)2}\cdots u_{\sigma(k)k}}\right]$
from (1.1).
The Weingarten function has been the subject of extensive research. This includes expansions using characters of symmetric groups, expressions through Jucys–Murphy elements, and series expansions as generating functions for monotone factorizations of permutations [Nov10, MN13]. In addition, a relation identity has been obtained between the Weingarten functions on $S_{k}$ and those on $S_{k-1}$ [CM17].
1.2. The problem of high degree integrals
A common shortoming of Weingarten calculus is that if one wants to evaluate an integral of degree strictly more than $n$ on the unitary or orthogonal group of dimension $n$, the Weingarten matrix is not invertible anymore. For details, we refer, for example, to [CMN22]. There exist fixes to this problem. The first one is to take the pseudo-inverse of the Weingarten matrix when it is not invertible. It is important to note that while a Weingarten formula exists irrespective of the degree, the Weingarten function is not unique when the degree becomes larger than the dimension.
We are aware of two additional routes to compute polynomial integrals of high degree against the Haar measure, and describe them briefly in what follows. Both workarounds are successful from a theoretical point of view in the sense that they allow to compute in principle the integrals. However, at this point we are unable to extract any useful information for concrete applications. The first attempt can be found as a remark of a paper of Collins and Śniady [CŚ06], and the authors, along with Fukuda, have a joint working project to expand on this topic, [CFM25]. Roughly, the idea is that any formal computation in the dimension, while in principle only well defined for a dimension that is large enough – and potentially larger than the dimension in which one wants to compute, doing the calculation formally, simplifying, and then evaluating in the target dimension will systematically give the correct result.
The second route was taken by Cioppa and Collins in the following unpbulished paper [CC13]. Here, the idea is to replace the indicator functions which are simple to compute but do not yield an orthogonal Gram matrix by much more complicated functions obtained from the Gelfand-Zetlin basis, i.e. exhibit matrix elements of the commutant of tensor representations of the unitary group.
The third approach is novel – it is actually the motivation for this paper. It comes from the following simple idea: from the point of view of differential geometry and measure theory, $\mathrm{U}(n)/\mathrm{U}(n-1)=S^{2n-1}$, and thus $\mathrm{U}(n)=S^{2n-1}\times\mathrm{U}(n-1)$ and this extends to the level of product of uniform measures. For a rigorous statement, we refer to Lemma 2.1. Therefore, we can understand the Haar measure at the level $n$ from the Haar measure at the level n1n-1 and the uniform measure on the sphere. But it turns out that the uniform measure on the sphere can be understood with arbitrary order directly thanks to a Gaussianization trick. This has been used in [BC08], and we also refer to the remark after Lemma 2.2 for more details.
This is the approach that we develop on this paper and it allows us to revisit the theory of Weingarten calculus from a completely different point of view and in principle solve completely the aforementioned problem. For that purpose, we need the concept of virtual isometries, which we describe in the forthcoming subsection.
1.3. Virtual isometries
Olshanski [Ols03] considered a space of virtual unitary matrices as projective limits of unitary groups. Additionally, Bourgade, Najnudel, and Nikeghbali [BNN13] independently explored this concept and referred to it as virtual isometry. In this paper, we will present an introduction based on the latter approach.
A virtual isometry is a sequence $(g_{n})_{n\geq 1}$ of unitary matrices such that for all $n\geq 1$, $g_{n}\in\mathrm{U}(n)$ and $\pi_{n}(g_{n})=g_{n-1}$. Here, a specific definition for the projection $\pi_{n}:\mathrm{U}(n)\to\mathrm{U}(n-1)$ is given by Neretin [Ner02] as follows. For a unitary matrix $g_{n}=(a_{ij})_{1\leq i,j\leq n}\in\mathrm{U}(n)$,
$\pi_{n}(g_{n})=\left(a_{ij}+\frac{a_{in}a_{nj}}{1-a_{nn}}\right)_{1\leq i,j\leq n-1}$
if $a_{nn}\not=1$, and $\pi_{n}(g_{n})=(a_{ij})_{1\leq i,j\leq n-1}$ if $a_{nn}=1$. Such virtual isometries serve as extensions of the virtual permutations first presented by Kerov, Olshanski, and Vershik [KOV93, KOV04].
Due to [BNN13] (see also [MNN20]), virtual isometries can be constructed inductively as follows: We equip $\mathbb{C}^{n}$ with the canonical scalar product: $\langle\bm{v},\bm{w}\rangle=\sum_{j=1}^{n}v_{j}\overline{w_{j}}$ for $\bm{v}=(v_{1},\dots,v_{n})^{\mathtt{T}}$, $\bm{w}=(w_{1},\dots,w_{n})^{\mathtt{T}}\in\mathbb{C}^{n}$. Here $A^{\mathtt{T}}$ and $\bm{v}^{\mathtt{T}}$ stand for the transpose of a matrix $A$ or a vector $\bm{v}$. Let $\bm{e}_{1},\dots,\bm{e}_{n}$ be the standard orthonormal basis of $\mathbb{C}^{n}$. Consider a sequence $(\bm{x}_{n})_{n\geq 1}$ of column vectors, where each $\bm{x}_{n}$ lies on the complex unit sphere of $\mathbb{C}^{n}$ for $n\geq 1$. We define a sequence of unitary matrices $r_{n}\in\mathrm{U}(n)$ in the following way:
  • $r_{n}$ is the identity matrix of size $n$ if $\bm{x}_{n}=\bm{e}_{n}$.
  • Otherwise, $r_{n}\in\mathrm{U}(n)$ is the unique complex reflection such that $r_{n}\bm{e}_{n}=\bm{x}_{n}$.
A virtual isometry $(g_{n})_{n\geq 1}$ is given by induction as follows:
$g_{1}:=r_{1}=\bm{x}_{1},\qquad g_{n}:=r_{n}\begin{pmatrix}g_{n-1}&\bm{0}\\ \bm{0}^{\mathtt{T}}&1\end{pmatrix}\quad\text{for $n\geq 2$}.$
One can check that $\pi_{n}(g_{n})=g_{n-1}$ in this construction.
The virtual isometry is consistent with Haar measure. Indeed, if $(X_{n})_{n\geq 1}$ is a sequence of independent random vectors, where each $X_{n}$ is uniformly distributed on the complex unit sphere of $\mathbb{C}^{n}$, and if we let $R_{n}$ and $U_{n}$ be the corresponding random matrices $r_{n}$ and $g_{n}$, respectively, in the above constructions, then $U_{n}$ follows the Haar measure on $\mathrm{U}(n)$.
A similar construction method for Haar measure is also provided by A. Hurwitz, refer to [DF17]. Virtual isometries have proven to be valuable in the analytic aspects of random matrix theory, enabling the attainment of almost sure results due to the correlations they provide between sequences of Haar-distributed unitary matrices, refer to [Ass22, CNN17].
1.4. Main results
Our goal is to provide a new perspective on Weingarten calculus based on the construction method of virtual isometry. The two main results of this paper are as follows.
The first main result is Theorem 2.4, where focus on a random complex reflection $R_{n}$ used in a construction involving virtual isometries. We establish Weingarten calculus for $R_{n}$, which provides a systematic method for computing the moments of the matrix elements of $R_{n}$. Instead of working directly with $R_{n}$, we will deal with $P_{n}:=I_{n}-R_{n}$ because $P_{n}$ is a rank-one matrix, making it easier to handle. Fortunately, the Weingarten calculus for $P_{n}=(p_{ij})_{1\leq i,j\leq n}$ has been settled in a nice form. In fact, the moment for the form
$\mathbb{E}\left[p_{i_{1}j_{1}}p_{i_{2}j_{2}}\cdots p_{i_{m}j_{m}}\overline{p_{i_{1}^{\prime}j_{1}^{\prime}}p_{i_{2}^{\prime}j_{2}^{\prime}}\cdots p_{i_{l}^{\prime}j_{l}^{\prime}}}\right]$
is explicitly given by a simple rational function in $n$. We provide many examples of calculating the moments of matrix elements of $R_{n}$ and $P_{n}$. For example, we obtain
$\mathbb{E}[|p_{11}p_{22}\cdots p_{kk}|^{2}]=\frac{2^{k}}{n^{2}(n+1)^{2}\cdots(n+k-1)^{2}}$
if $k<n$. From the construction of the virtual isometry, the moments for $P_{n}$ essentially contain all the information of a Haar-distributed unitary matrix $U_{n}$. Therefore, this allows for the calculation of the moments of Haar-distributed unitary matrices in a way that is different from Weingarten calculus theoretically.
Our second main result is Theorem 3.3, in which we give an explicit relationship is established between the unitary Weingarten functions of dimensions $n$ and n1n-1. More especially, if $n>k$, the Weingarten function $\operatorname{Wg}_{k,n}$ on the symmetric group $S_{k}$ is expressed as a convolution product in the group algebra of $S_{k}$:
$\operatorname{Wg}_{k,n}=\operatorname{as}_{k,n}*\operatorname{Wg}_{k,n-1}$
where $\operatorname{as}_{k,n}$ is a new class function on $S_{k}$ and $\operatorname{Wg}_{k,n-1}$ is the Weingarten function of lower dimension n1n-1. We named the function $\operatorname{as}_{k,n}$ an ascension function and investigated its basic properties. For example, finding the inverse of function $\operatorname{as}_{k,n}$ with respect to the convolution product in the symmetric group algebra is straightforward. The assumption $n>k$ is necessary in the above formula. Although a similar expression can be derived for the case when $n=k$ with minor adjustments, it is important to emphasize that the resulting expression does not apply universally for all values of $n$ and $k$. Consequently, we have obtained a new representation of the Weingarten function utilizing the ascension functions and related constructs (see Proposition 3.11).
These two main results are discussed in Sections 2 and 3, respectively. Importantly, let us add that Section 2.5 explains how to use these results systematically to compute integrals of polynomials of arbitrary degree, thus providing a new approach to the problem mentioned in Section 1.2.
Extending these results to the orthogonal group is not particularly challenging; This will be presented at a later opportunity.
At the end of the introduction, we comment on symbols used. Throughout this article, we use the notation
$a^{\uparrow k}=\prod_{j=0}^{k-1}(a+j)$
for a complex number $a$ and a positive integer $k$. For convenience, we also set $a^{\uparrow 0}=1$. For clarity, we will use both $\delta_{ij}$ and $\delta(i,j)$ as Kronecker’s delta. The symbol $I_{n}$ represents the identity matrix of size $n$.
Acknowledgments
B. C. is supported by JSPS Grant-in-Aid Scientific Research (A) no. 25H00593, and Challenging Research (Exploratory) no. 20K20882 and 23K17299.
S. M. is supported by JSPS Grant-in-Aid Scientific Research (C) no. 22K03233.
The idea to work on this project arose from discussions of B. C. with Sasha Bufetov and a talk by Ashkan Nikeghbali at the June 2023 RIMS conference on “Random Matrices and Applications”. We acknowledge discussions with them and the support of RIMS.
2. Weingarten calculus for complex reflections
2.1. Complex reflection in the construction of virtual isometry
We equip $\mathbb{C}^{n}$ with the canonical scalar product: $\langle\bm{v},\bm{w}\rangle=\sum_{j=1}^{n}v_{j}\overline{w_{j}}$ for $\bm{v}=(v_{1},\dots,v_{n})^{\mathtt{T}}$, $\bm{w}=(w_{1},\dots,w_{n})^{\mathtt{T}}\in\mathbb{C}^{n}$. Let $\bm{e}_{1},\dots,\bm{e}_{n}$ be the standard orthonormal basis of $\mathbb{C}^{n}$.
A complex reflection on $\mathbb{C}^{n}$ is a unitary transformation $\mathcal{R}$ such that it is either the identity map $\mathrm{Id}$ or the rank of $\mathrm{Id}$-$\mathcal{R}$ is $1$. Every complex reflection can be represented as
$\mathcal{R}_{\bm{a},\alpha}(\bm{v})=\bm{v}-(1-\alpha)\frac{\langle\bm{v},\bm{a}\rangle}{\langle\bm{a},\bm{a}\rangle}\bm{a}\qquad(\bm{v}\in\mathbb{C}^{n}),$
where $\bm{a}$ is some nonzero vector and $\alpha$ is a complex number with $|\alpha|=1$. Observe that $\mathcal{R}_{\bm{a},\alpha}(\bm{a})=\alpha\bm{a}$ and $\mathcal{R}_{\bm{a},\alpha}(\bm{w})=\bm{w}$ if $\bm{w}$ is orthogonal to $\bm{a}$. Given two distinct unit vectors $\bm{e}$ and $\bm{m}$, there exists a unique reflection $\mathcal{R}$ such that $\mathcal{R}(\bm{e})=\bm{m}$, and it is given by $\mathcal{R}_{\bm{m}-\bm{e},\alpha}$ with $\alpha=-\frac{1-\langle\bm{m},\bm{e}\rangle}{1-\langle\bm{e},\bm{m}\rangle}$.
Let
$X=X_{n}=(x_{1},x_{2},\dots,x_{n})^{\mathtt{T}}$
be a uniform random vector on the unit sphere of $\mathbb{C}^{n}$. Since $X$ is almost surely different from $\bm{e}_{n}$, there exists a unique reflection $\mathcal{R}$ such that $\mathcal{R}(\bm{e}_{n})=X$ as described above. We denote by $R=R_{n}=(r_{ij})_{1\leq i,j\leq n}$ its matrix representation with respect to the basis $\bm{e}_{1},\dots,\bm{e}_{n}$. It is straightforward to see that
[…]
In particular, the last column of $R$ is $X_{n}$.
The following lemma is the key to our approach. For a proof, we refer to [BNN13, Proposition 3.1] or [BNR11, Theorem 1]. See also [BHNY08, Proposition 2.1].
Lemma 2.1.
Let $n\geq 2$. Let $V$ be a Haar-distributed unitary matrix from $\mathrm{U}(n-1)$ and set
(2.2) $\tilde{V}:=V\oplus(1)=\begin{pmatrix}V&\bm{0}\\ \bm{0}&1\end{pmatrix}\in\mathrm{U}(n).$
Let $X=(x_{1},x_{2},\dots,x_{n})^{\mathtt{T}}$ be a uniform random vector on the unit sphere of $\mathbb{C}^{n}$, independent of $V$, and we construct the random unitary matrix $R=(r_{ij})$ as described in (2.1). Then the product $U:=R\tilde{V}$ follows the Haar measure on $\mathrm{U}(n)$.
The next lemma will be used several times.
Lemma 2.2.
Let $X=(x_{1},\dots,x_{n})^{\mathtt{T}}$ be a uniform vector on the unit sphere of $\mathbb{C}^{n}$. For $2n$ nonnegative integers $m_{1},\dots,m_{n},l_{1},\dots,l_{n}$, the moment
$\mathbb{E}[x_{1}^{m_{1}}x_{2}^{m_{2}}\cdots x_{n}^{m_{n}}\overline{x_{1}^{l_{1}}x_{2}^{l_{2}}\cdots x_{n}^{l_{n}}}]$
survives only if $m_{j}=l_{j}$ for all $1\leq j\leq n$. In this case, we have
$\mathbb{E}[|x_{1}|^{2m_{1}}|x_{2}|^{2m_{2}}\cdots|x_{n}|^{2m_{n}}]=\frac{m_{1}!\,m_{2}!\,\cdots\,m_{n}!}{n^{\uparrow m}},$
where $m=m_{1}+m_{2}+\cdots+m_{n}$.
Proof.
The random vector $X$ has the same distribution as the first column of a Haar-distributed unitary matrix from $\mathrm{U}(n)$. This lemma is an easy example of Weingarten calculus. In fact, using (1.1), we can see that
$\mathbb{E}[|x_{1}|^{2m_{1}}|x_{2}|^{2m_{2}}\cdots|x_{n}|^{2m_{n}}]$ $=m_{1}!\,m_{2}!\,\cdots\,m_{n}!\sum_{\sigma\in S_{m}}\operatorname{Wg}_{m,n}(\sigma)$
$=\frac{m_{1}!\,m_{2}!\,\cdots\,m_{n}!}{n(n+1)\cdots(n+m-1)}.$
The second equality can be easily proved, for example, by using the character expansion (3.2) given later. ∎
Remark 2.3.
This lemma can also be obtained through a trick that is sometimes called the gaussianization method. Indeed, let $y_{1},\ldots,y_{n}$ be $n$ i.i.d standard complex gaussian variables. Then, the random vector $y/\|y\|_{2}$ has the same distribution as the uniformly distributed vector $x=(x_{1},\ldots,x_{n})$ on the norm $1$ sphere. Conversely, if $y$ and $x$ are independent, $\|y\|_{2}\cdot x$ has the same distribution as $y$. Using the complex version of Wick’s theorem and the fact that $\|y\|$ has a complex Chi-square distribution, thus satisfying $\mathbb{E}(\|y\|_{2}^{2k})=n(n+1)\cdots(n+k-1)$, the proof follows. For applications and related materials, we refer to Theorem 6.1 of [BC08]. We also refer to [DE02], in which a Chi-square distribution also appears in a tridiagonalization of a related random matrix.
2.2. Weingarten calculus for a rank-one matrix
Instead of working directly with $R$, We deal with $P=(p_{ij})_{1\leq i,j\leq n}:=I-R$ because $P$ is a rank-one matrix, thus easier to handle. By (2.1) we have
[…]
Let us fix some notation. For a sequence $\bm{i}=(i_{1},\dots,i_{m})$ of positive integers and a positive integer $k$, we denote by $\mathsf{m}_{k}(\bm{i})$ the multiplicity of $k$ in $\bm{i}$:
$\mathsf{m}_{k}(\bm{i})=|\{h\in\{1,2,\dots,m\}\ |\ i_{h}=k\}|.$
For two sequences $\bm{i}=(i_{1},\dots,i_{m})$ and $\bm{j}=(j_{1},\dots,j_{l})$, we define a combined sequence $\bm{i}\sqcup\bm{j}=(i_{1},\dots,i_{m},j_{1},\dots,j_{l})$. In what follows, the order of the letters is not significant. Note that $\mathsf{m}_{k}(\bm{i}\sqcup\bm{j})=\mathsf{m}_{k}(\bm{i})+\mathsf{m}_{k}(\bm{j})$.
The following theorem is our first main result.
Theorem 2.4.
Let $m,l$ be nonnegative integers, and consider four sequences
$\bm{i}=(i_{1},\dots,i_{m}),\quad\bm{j}=(j_{1},\dots,j_{m}),\quad\bm{i}^{\prime}=(i_{1}^{\prime},\dots,i_{l}^{\prime}),\quad\bm{j}^{\prime}=(j_{1}^{\prime},\dots,j_{l}^{\prime}).$
of positive integers in $\{1,\dots,n\}$. Then the moment
$\mathbb{E}\left[p_{i_{1}j_{1}}\cdots p_{i_{m}j_{m}}\overline{p_{i_{1}^{\prime}j_{1}^{\prime}}\cdots p_{i_{l}^{\prime}j_{l}^{\prime}}}\right]$
vanishes unless $\bm{i}\sqcup\bm{j}^{\prime}$ is a permutation of $\bm{j}\sqcup\bm{i}^{\prime}$. In this case, the moment is equal to
$\alpha_{1}!\cdots\alpha_{n-1}!\frac{n^{\uparrow\alpha_{n}}}{n^{\uparrow m}n^{\uparrow l}},$
where
$\alpha_{k}:=\mathsf{m}_{k}(\bm{i}\sqcup\bm{j}^{\prime})=\mathsf{m}_{k}(\bm{j}\sqcup\bm{i}^{\prime})\qquad(k=1,\dots,n).$
We postpone the proof of Theorem 2.4 to the next subsection.
Example 2.5.
  1. (1)
    Suppose $n>1$ and consider
    $\mathbb{E}[|p_{11}|^{4}|p_{nn}|^{2}]=\mathbb{E}[(p_{11})^{2}p_{nn}(\overline{p_{11}})^{2}\overline{p_{nn}}].$
    Apply Theorem 2.4 with
    $\bm{i}=\bm{j}=\bm{i}^{\prime}=\bm{j}^{\prime}=(1,1,n).$
    Then
    $\mathbb{E}[|p_{11}|^{4}|p_{nn}|^{2}]=4!\frac{n^{\uparrow 2}}{n^{\uparrow 3}n^{\uparrow 3}}=\frac{24}{n(n+1)(n+2)^{2}}.$
  2. (2)
    Suppose $n>2$ and consider $\mathbb{E}[(p_{12})^{2}(p_{n1})^{2}(p_{nn})^{3}(\overline{p_{n2}})^{2}]$. The data is
    $\bm{i}=(1,1,n,n,n,n,n),\quad\bm{j}=(2,2,1,1,n,n,n),\quad\bm{i}^{\prime}=(n,n),\quad\bm{j}^{\prime}=(2,2).$
    Then
    $\mathbb{E}[(p_{12})^{2}(p_{n1})^{2}(p_{nn})^{3}(\overline{p_{n2}})^{2}]=2!\,2!\,\frac{n^{\uparrow 5}}{n^{\uparrow 7}n^{\uparrow 2}}=\frac{4}{n(n+1)(n+5)(n+6)}.$
  3. (3)
    Suppose $n>3$ and consider $\mathbb{E}[p_{12}p_{21}(p_{nn})^{4}(\overline{p_{33}})^{2}(\overline{p_{nn}})^{3}]$. The data is
    $\bm{i}=(1,2,n,n,n,n),\quad\bm{j}=(2,1,n,n,n,n),\quad\bm{i}^{\prime}=\bm{j}^{\prime}=(3,3,n,n,n).$
    Then
    $\mathbb{E}[p_{12}p_{21}(p_{nn})^{4}(\overline{p_{33}})^{2}(\overline{p_{nn}})^{3}]$ $=1!\,1!\,2!\,\frac{n^{\uparrow 7}}{n^{\uparrow 6}n^{\uparrow 5}}$
    $=\frac{2(n+6)}{n(n+1)(n+2)(n+3)(n+4)}.$
2.3. Proof of the theorem
Let us rewrite Theorem 2.4 in another form. For an $n\times n$ matrix $A=(a_{ij})_{1\leq i,j\leq n}$ and each $1\leq k\leq n$, we denote by $a_{k\bullet}$ the sum of all entries in the $k$-th row of $A$ and by $a_{\bullet k}$ the sum of all entries in the $k$-th column of $A$:
$a_{k\bullet}=\sum_{j=1}^{n}a_{kj},\qquad a_{\bullet k}=\sum_{i=1}^{n}a_{ik}.$
Furthermore, we define
$\Sigma_{A}=\sum_{i=1}^{n}\sum_{j=1}^{n}a_{ij}=\sum_{k=1}^{n}a_{k\bullet}=\sum_{k=1}^{n}a_{\bullet k}.$
It is immediate to see that Theorem 2.4 is equivalent to Theorem 2.6.
Theorem 2.6.
Let $A=(a_{ij})_{1\leq i,j\leq n}$ and $B=(b_{ij})_{1\leq i,j\leq n}$ be two matrices whose entries are nonnegative integers. Then the moment
$\mathbb{E}\left[\prod_{i,j=1}^{n}(p_{ij})^{a_{ij}}(\overline{p_{ij}})^{b_{ij}}\right]$
vanishes unless
$a_{\bullet k}+b_{k\bullet}=a_{k\bullet}+b_{\bullet k}=:\alpha_{k}$
for all $k=1,2,\dots,n$. In this case, the moment is equal to
$\alpha_{1}!\cdots\alpha_{n-1}!\frac{n^{\uparrow\alpha_{n}}}{n^{\uparrow\Sigma_{A}}n^{\uparrow\Sigma_{B}}}.$
Proof.
By the definition (2.3) of $(p_{ij})$, we have
$\prod_{i,j=1}^{n}(p_{ij})^{a_{ij}}(\overline{p_{ij}})^{b_{ij}}=(-1)^{c}\prod_{k=1}^{n-1}\left(x_{k}^{a_{\bullet k}+b_{k\bullet}}\overline{x_{k}}^{a_{k\bullet}+b_{\bullet k}}\right)\cdot(1-x_{n})^{s}(1-\overline{x_{n}})^{t},$
where
$c$ $=\sum_{j=1}^{n-1}(a_{nj}+b_{nj})+\sum_{i=1}^{n-1}(a_{in}+b_{in}),$
$s$ $=a_{n\bullet}-\sum_{j=1}^{n-1}b_{\bullet j},\qquad t=b_{n\bullet}-\sum_{j=1}^{n-1}a_{\bullet j}.$
Since $|x_{n}|<1$ almost surely, the generalized binomial theorem implies that
$\mathbb{E}\left[\prod_{i,j=1}^{n}(p_{ij})^{a_{ij}}(\overline{p_{ij}})^{b_{ij}}\right]$
$=(-1)^{c}\sum_{m=0}^{\infty}\sum_{l=0}^{\infty}\frac{(-s)^{\uparrow m}(-t)^{\uparrow l}}{m!\,l!}\mathbb{E}\left[\prod_{k=1}^{n-1}\left(x_{k}^{a_{\bullet k}+b_{k\bullet}}\overline{x_{k}}^{a_{k\bullet}+b_{\bullet k}}\right)\cdot x_{n}^{m}\overline{x_{n}}^{l}\right].$
Since we are integrating on the unit sphere, the interchange of the infinite sum and the expectation follows from Fubini’s theorem. Applying Lemma 2.2 to each term, it survives only if equalities $a_{\bullet k}+b_{k\bullet}=a_{k\bullet}+b_{\bullet k}$ hold for all $k=1,2,\dots,n-1$, and if $m=l$ holds. In this case, we also have $a_{\bullet n}+b_{n\bullet}=a_{n\bullet}+b_{\bullet n}$ and $c\equiv 0\pmod{2}$, which leads to the following equations:
(2.4) $s+\alpha_{1}+\cdots+\alpha_{n-1}$ $=\Sigma_{A},$
$t+\alpha_{1}+\cdots+\alpha_{n-1}$ $=\Sigma_{B},$
$s+t+\alpha_{1}+\cdots+\alpha_{n-1}$ $=\alpha_{n}.$
Furthermore, from Lemma 2.2 again, we have
$\mathbb{E}\left[\prod_{i,j=1}^{n}(p_{ij})^{a_{ij}}(\overline{p_{ij}})^{b_{ij}}\right]=\alpha_{1}!\cdots\alpha_{n-1}!\sum_{m=0}^{\infty}\frac{(-s)^{\uparrow m}(-t)^{\uparrow m}}{m!\,n^{\uparrow(\alpha_{1}+\cdots+\alpha_{n-1}+m)}}.$
We continue the calculations of the last series.
$\sum_{m=0}^{\infty}\frac{(-s)^{\uparrow m}(-t)^{\uparrow m}}{m!\,n^{\uparrow(\alpha_{1}+\cdots+\alpha_{n-1}+m)}}$
$=\frac{1}{n^{\uparrow(\alpha_{1}+\cdots+\alpha_{n-1})}}\sum_{m=0}^{\infty}\frac{(-s)^{\uparrow m}(-t)^{\uparrow m}}{m!\,(n+\alpha_{1}+\cdots+\alpha_{n-1})^{\uparrow m}}$
$=\frac{1}{n^{\uparrow(\alpha_{1}+\cdots+\alpha_{n-1})}}\cdot F(-s,-t;n+\alpha_{1}+\cdots+\alpha_{n-1};1),$
where $F(\alpha,\beta;\gamma;x)$ is the Gauss hypergeometric series. The special value $F(\alpha,\beta;\gamma;1)$ makes sense if $\gamma-(\alpha+\beta)>0$, which is fulfilled in our setting. Using the well-known formula $F(\alpha,\beta;\gamma;1)=\frac{\Gamma(\gamma)\Gamma(\gamma-\alpha-\beta)}{\Gamma(\gamma-\alpha)\Gamma(\gamma-\beta)}$ (see, e.g., [AAR00, Theorem 2.2.2]), where $\Gamma(x)$ is the gamma function, we obtain
$F(-s,-t;n+\alpha_{1}+\cdots+\alpha_{n-1};1)=\frac{\Gamma(n+\alpha_{1}+\cdots+\alpha_{n-1})\Gamma(n+\alpha_{n})}{\Gamma(n+\Sigma_{A})\Gamma(n+\Sigma_{B})},$
where we used (2.4). The result follows by using $n^{\uparrow k}=\frac{\Gamma(n+k)}{\Gamma(n)}$. ∎
2.4. Some examples for reflection matrices
By utilizing Theorem 2.4 or Theorem 2.6, we can theoretically compute the moments for $(r_{ij})$. However, when there are many diagonal elements $r_{ii}$, there are many factors of the form $(1-p_{ii})$, which complicates the calculations using the theorems. While it seems complex in general, some calculations can be performed to a certain extent as follows.
Proposition 2.7.
Let $m,l,q$ be nonnegative integers, and consider four sequences
$\bm{i}=(i_{1},\dots,i_{m}),\quad\bm{j}=(j_{1},\dots,j_{m}),\quad\bm{i}^{\prime}=(i_{1}^{\prime},\dots,i_{l}^{\prime}),\quad\bm{j}^{\prime}=(j_{1}^{\prime},\dots,j_{l}^{\prime}).$
Then, for each $1\leq s\leq n$ the moment
$\mathbb{E}\left[p_{i_{1}j_{1}}\cdots p_{i_{m}j_{m}}\overline{p_{i_{1}^{\prime}j_{1}^{\prime}}\cdots p_{i_{l}^{\prime}j_{l}^{\prime}}}r_{ss}^{q}\right]$
survives only if $\bm{i}\sqcup\bm{j}^{\prime}$ is a permutation of $\bm{j}\sqcup\bm{i}^{\prime}$. In this case, the moment is equal to
$\alpha_{1}!\cdots\alpha_{n-1}!\,\frac{n^{\uparrow\alpha_{n}}}{n^{\uparrow(m+q)}n^{\uparrow l}}\times\begin{cases}(n+m-\alpha_{s}-1)^{\uparrow q}&\text{if $s<n$};\\ (m-\alpha_{n})^{\uparrow q}&\text{if $s=n$},\end{cases}$
where $\alpha_{k}:=\mathsf{m}_{k}(\bm{i}\sqcup\bm{j}^{\prime})=\mathsf{m}_{k}(\bm{j}\sqcup\bm{i}^{\prime})$ for all $k=1,2,\dots,n$. Here the last factor may be zero.
Proof.
By the binomial theorem with $r_{ss}^{q}=(1-p_{ss})^{q}$, we have
$\mathbb{E}\left[p_{i_{1}j_{1}}\cdots p_{i_{m}j_{m}}\overline{p_{i_{1}^{\prime}j_{1}^{\prime}}\cdots p_{i_{l}^{\prime}j_{l}^{\prime}}}r_{ss}^{q}\right]$
(2.5) $=\sum_{k=0}^{q}\frac{(-q)^{\uparrow k}}{k!}\mathbb{E}\left[p_{i_{1}j_{1}}\cdots p_{i_{m}j_{m}}\overline{p_{i_{1}^{\prime}j_{1}^{\prime}}\cdots p_{i_{l}^{\prime}j_{l}^{\prime}}}p_{ss}^{k}\right].$
First, we deal with the case $s<n$. Applying Theorem 2.4 to each term, we have
(2.5) $=\sum_{k=0}^{q}\frac{(-q)^{\uparrow k}}{k!}\alpha_{1}!\,\cdots\,\alpha_{n-1}!\cdot\frac{(\alpha_{s}+k)!}{\alpha_{s}!}\cdot\frac{n^{\uparrow\alpha_{n}}}{n^{\uparrow(m+k)}n^{\uparrow l}}$
$=\alpha_{1}!\,\cdots\,\alpha_{n-1}!\cdot\frac{n^{\uparrow\alpha_{n}}}{n^{\uparrow m}n^{\uparrow l}}\sum_{k=0}^{q}\frac{(-q)^{\uparrow k}(\alpha_{s}+1)^{\uparrow k}}{k!\,(n+m)^{\uparrow k}}.$
Using a similar method to the proof of Theorem 2.4, we view the last sum as the Gauss hypergeometric series, and obtain
$=\alpha_{1}!\,\cdots\,\alpha_{n-1}!\cdot\frac{n^{\uparrow\alpha_{n}}}{n^{\uparrow m}n^{\uparrow l}}\cdot\frac{(n+m-\alpha_{s}-1)^{\uparrow q}}{(n+m)^{\uparrow q}}$
$=\alpha_{1}!\,\cdots\,\alpha_{n-1}!\cdot\frac{n^{\uparrow\alpha_{n}}}{n^{\uparrow(m+q)}n^{\uparrow l}}\cdot(n+m-\alpha_{s}-1)^{\uparrow q}.$
This concludes the proof for the case when $s<n$.
Next, we deal with the case $s=n$. A similar approach gives
(2.5) $=\sum_{k=0}^{q}\frac{(-q)^{\uparrow k}}{k!}\alpha_{1}!\cdots\alpha_{n-1}!\,\frac{n^{\uparrow(\alpha_{n}+k)}}{n^{\uparrow(m+k)}n^{\uparrow l}}$
$=\alpha_{1}!\cdots\alpha_{n-1}!\,\frac{n^{\uparrow\alpha_{n}}}{n^{\uparrow m}n^{\uparrow l}}\sum_{k=0}^{q}\frac{(-q)^{\uparrow k}(n+\alpha_{n})^{\uparrow k}}{k!\,(n+m)^{\uparrow k}}$
$=\alpha_{1}!\cdots\alpha_{n-1}!\,\frac{n^{\uparrow\alpha_{n}}}{n^{\uparrow m}n^{\uparrow l}}\cdot\frac{(m-\alpha_{n})^{\uparrow q}}{(n+m)^{\uparrow q}}.$
This concludes the proof for the case when $s=n$. ∎
Example 2.8.
Let $q$ be a positive integer.
$\mathbb{E}\left[r_{ss}^{q}\right]=\begin{cases}\frac{n-1}{n+q-1}&\text{if $1\leq s<n$};\\ 0&\text{if $s=n$}.\end{cases}$
Example 2.9.
If $n>2$, then
$\mathbb{E}\left[r_{12}^{2}r_{n1}^{2}\overline{r_{n2}}^{2}r_{22}^{3}\right]=\mathbb{E}\left[p_{12}^{2}p_{n1}^{2}\overline{p_{n2}}^{2}r_{22}^{3}\right]$ $=\frac{4}{n(n+4)(n+5)(n+6)},$
$\mathbb{E}\left[r_{12}^{2}r_{n1}^{2}\overline{r_{n2}}^{2}r_{nn}^{3}\right]=\mathbb{E}\left[p_{12}^{2}p_{n1}^{2}\overline{p_{n2}}^{2}r_{nn}^{3}\right]$ $=\frac{96}{n^{\uparrow 7}}.$
Proposition 2.10.
Let $m,l$ be nonnegative integers, and consider four sequences
$\bm{i}=(i_{1},\dots,i_{m}),\quad\bm{j}=(j_{1},\dots,j_{m}),\quad\bm{i}^{\prime}=(i_{1}^{\prime},\dots,i_{l}^{\prime}),\quad\bm{j}^{\prime}=(j_{1}^{\prime},\dots,j_{l}^{\prime}).$
Then, for each $1\leq s\leq n$, the moment
$\mathbb{E}\left[p_{i_{1}j_{1}}\cdots p_{i_{m}j_{m}}\overline{p_{i_{1}^{\prime}j_{1}^{\prime}}\cdots p_{i_{l}^{\prime}j_{l}^{\prime}}}|r_{ss}|^{2}\right]$
survives only if $\bm{i}\sqcup\bm{j}^{\prime}$ is a permutation of $\bm{j}\sqcup\bm{i}^{\prime}$. In this case, the moment is equal to
$\alpha_{1}!\cdots\alpha_{n-1}!\,\frac{n^{\uparrow\alpha_{n}}}{n^{\uparrow(m+1)}n^{\uparrow(l+1)}}$
$\times\begin{cases}\Big((n+m)(n+l)-(\alpha_{s}+1)(2n+m+l)+(\alpha_{s}+1)(\alpha_{s}+2)\Big)&\text{if $s<n$};\\ \Big(n+(m-\alpha_{n})(l-\alpha_{n})+\alpha_{n}\Big)&\text{if $s=n$},\end{cases}$
where $\alpha_{k}:=\mathsf{m}_{k}(\bm{i}\sqcup\bm{j}^{\prime})=\mathsf{m}_{k}(\bm{j}\sqcup\bm{i}^{\prime})$ for all $k=1,2,\dots,n$.
Proof.
Expand $|r_{ss}|^{2}=1-p_{ss}-\overline{p_{ss}}+p_{ss}\overline{p_{ss}}$ and apply Theorem 2.4. We omit the detailed calculations. ∎
Example 2.11.
Let $n>1$. From Proposition 2.10 we have
$\mathbb{E}[|r_{nn}|^{2}]=\frac{1}{n},\quad\mathbb{E}[p_{11}|r_{nn}|^{2}]=\frac{1}{n(n+1)},\quad\mathbb{E}[p_{11}\overline{p_{11}}|r_{nn}|^{2}]=\frac{2}{n^{2}(n+1)}.$
Using these, we obtain the following:
$\mathbb{E}[|r_{11}r_{nn}|^{2}]$ $=\mathbb{E}[(1-p_{11})(1-\overline{p_{11}})|r_{nn}|^{2}]$
$=\frac{1}{n}-2\cdot\frac{1}{n(n+1)}+\frac{2}{n^{2}(n+1)}$
$=\frac{n^{2}-n+2}{n^{2}(n+1)}.$
Proposition 2.12.
Let $m$ be a positive integer, and consider two sequences
$\bm{i}=(i_{1},\dots,i_{m}),\quad\bm{j}=(j_{1},\dots,j_{m}).$
Assume that the components of $\bm{i}$ are distinct and that $\bm{i}$ does not contain $n$. Then the moment
$\mathbb{E}\left[r_{i_{1}j_{1}}r_{i_{2}j_{2}}\cdots r_{i_{m}j_{m}}\right]$
vanishes unless $\bm{j}$ is a permutation of $\bm{i}$. In this case, the moment is equal to
(2.6) $\sum_{t=0}^{f}(-1)^{m-t}\binom{f}{t}\frac{1}{n^{\uparrow(m-t)}},$
where $f=f(\bm{i},\bm{j})=|\{h\in\{1,2,\dots,m\}|\ i_{h}=j_{h}\}|$.
Proof.
We may suppose that $i_{h}=j_{h}$ for $1\leq h\leq f$ and $i_{h}\not=j_{h}$ for $f+1\leq h\leq m$. We define:
$\bm{f}:=(i_{1},i_{2},\dots,i_{f}),$
$\bm{i}_{0}:=(i_{f+1},i_{f+2},\dots,i_{m}),\quad\bm{j}_{0}:=(j_{f+1},j_{f+2},\dots,j_{m}).$
The sequence $\bm{f}$ can be identified with the set $\{i_{1},i_{2},\dots,i_{f}\}$ since its components are distinct. We then have:
$\mathbb{E}\left[r_{i_{1}j_{1}}r_{i_{2}j_{2}}\cdots r_{i_{m}j_{m}}\right]$ $=\mathbb{E}\left[\prod_{i\in\bm{f}}(1-p_{i,i})\cdot\prod_{h=f+1}^{m}(-p_{i_{h},j_{h}})\right]$
$=\sum_{\bm{k}:\bm{k}\subset\bm{f}}\mathbb{E}\left[\prod_{i\in\bm{k}}(-p_{i,i})\cdot\prod_{h=f+1}^{m}(-p_{i_{h},j_{h}})\right],$
where $\bm{k}$ runs over all subsets of $\bm{f}$. Applying Theorem 2.4 to each term, we find that a term corresponding to $\bm{k}$ survives only if $\bm{k}\sqcup\bm{j}_{0}$ is a permutation of $\bm{k}\sqcup\bm{i}_{0}$. In other words, $\bm{j}$ must be a permutation of $\bm{i}$. Furthermore, we can continue the calculation as follows:
$\mathbb{E}\left[r_{i_{1}j_{1}}r_{i_{2}j_{2}}\cdots r_{i_{m}j_{m}}\right]=\sum_{s=0}^{f}\binom{f}{s}(-1)^{s+m-f}\frac{1}{n^{\uparrow(s+m-f)}}.$
The desired form is obtained by changing $s\mapsto t=f-s$. ∎
The following example plays a very important role from a theoretical point of view in Section 3.
Example 2.13.
For $k<n$ and $\sigma\in S_{k}$, we have
$\mathbb{E}[r_{1\sigma(1)}\cdots r_{k\sigma(k)}]=\sum_{t=0}^{\mathsf{f}(\sigma)}(-1)^{k-t}\binom{\mathsf{f}(\sigma)}{t}\frac{1}{n^{\uparrow(k-t)}},$
where $\mathsf{f}(\sigma)$ is the number of fixed points in $\sigma$:
$\mathsf{f}(\sigma)=|\{i\in\{1,2,\dots,k\}\ |\ \sigma(i)=i\}|.$
Remark 2.14.
We can express the function given in (2.6) by using Kummer’s confluent hypergeometric function
$M(a,b,z)=\sum_{k=0}^{\infty}\frac{a^{\uparrow k}}{b^{\uparrow k}}\frac{z^{k}}{k!}.$
Indeed,
$\sum_{t=0}^{f}(-1)^{m-t}\binom{f}{t}\frac{1}{n^{\uparrow(m-t)}}=\frac{(-1)^{m-f}}{n^{\uparrow(m-f)}}M(-f,n+m-f,1).$
2.5. Calculation of Haar-distributed matrices from reflection matrices
The following lemma is a reformulation of Lemma 2.1.
Lemma 2.15.
Let $X_{1},X_{2},\dots,X_{n}$ be independent random vectors, where each $X_{k}$ is uniformly distributed on the unit sphere of $\mathbb{C}^{k}$. For each $1\leq k\leq n$, we denote by $R_{k}$ the corresponding reflection matrix of size $k$ given in a manner similar to (2.1), and we define an $n\times n$ matrix $\tilde{R}_{k}=R_{k}\oplus I_{n-k}$. Then, a Haar-distributed unitary matrix $U$ on $\mathrm{U}(n)$ has the same distribution as the product $\tilde{R}_{n}\tilde{R}_{n-1}\cdots\tilde{R}_{1}$. In particular, the last $k$ columns of $U$ match those of $\tilde{R}_{n}\tilde{R}_{n-1}\cdots\tilde{R}_{n-k+1}$.
Thanks to this lemma, it is possible to calculate moments of polynomial functions on the unitary group by a method different from the Weingarten calculus, using Theorem 2.4. In this subsection, we provide some examples of such calculations, that indicate that the approach described in this paper allows to recover Weingarten functions.
Let $n\geq 2$, and let $U=(u_{ij})$ be a Haar-distributed unitary matrix on $\mathrm{U}(n)$. From Lemma 2.15, the last two columns of $U$ are expressed as follows:
$u_{i,n-1}=\sum_{j=1}^{n-1}r_{i,j}s_{j,n-1},\qquad u_{i,n}=r_{i,n}\qquad(1\leq i\leq n),$
where $r_{i,j}$ are matrix elements of $R_{n}$ and $s_{i,j}$ are those of $R_{n-1}$. Let us calculate the following moment.
$\mathbb{E}\left[u_{n-1,n-1}u_{n,n}\overline{u_{n-1,n}u_{n,n-1}}\right].$
Using the Weingaten calculus (1.1), it is immediately clear that this value is equal to $\operatorname{Wg}_{2,n}((1\ 2))=-\frac{1}{n(n+1)(n-1)}$, but here we shall deliberately take a longer route to calculate it. Since $(r_{ij})$ and $(s_{ij})$ are independent, we have
$\mathbb{E}\left[u_{n-1,n-1}u_{n,n}\overline{u_{n-1,n}u_{n,n-1}}\right]$
$=\sum_{j=1}^{n-1}\sum_{l=1}^{n-1}\mathbb{E}[r_{n-1,j}r_{n,n}\overline{r_{n-1,n}r_{n,l}}]\mathbb{E}[s_{j,n-1}\overline{s_{l,n-1}}].$
Example 2.11 or Proposition 2.7 imply that
$\mathbb{E}[s_{j,n-1}\overline{s_{l,n-1}}]$ $=\frac{\delta_{j,l}}{n-1}$ $\text{for $j,l\leq n-1$};$
$\mathbb{E}[r_{n-1,j}r_{n,n}\overline{r_{n-1,n}r_{n,j}}]$ $=0$ $\text{for $j<n-1$},$
and
$\mathbb{E}[r_{n-1,n-1}r_{n,n}\overline{r_{n-1,n}r_{n,n-1}}]$ $=\mathbb{E}[(1-p_{n-1,n-1})\overline{p_{n-1,n}p_{n,n-1}}r_{n,n}]$
$=-\frac{1}{n(n+1)}.$
Thus, we have obtained
$\mathbb{E}\left[u_{n-1,n-1}u_{n,n}\overline{u_{n-1,n}u_{n,n-1}}\right]$
$=\mathbb{E}[r_{n-1,n-1}r_{n,n}\overline{r_{n-1,n}r_{n,n-1}}]\mathbb{E}[s_{n-1,n-1}\overline{s_{n-1,n-1}}]$
$=-\frac{1}{n(n+1)(n-1)}$
as desired.
Let us perform a similar calculation for $\mathbb{E}\left[u_{n-1,n-1}u_{n,n}\overline{u_{n-1,n-1}u_{n,n}}\right]$, which equals $\operatorname{Wg}_{2,n}(e_{2})=\frac{1}{(n-1)(n+1)}$ by Weingarten calculus. We see that:
$\mathbb{E}\left[u_{n-1,n-1}u_{n,n}\overline{u_{n-1,n-1}u_{n,n}}\right]$ $=\sum_{j=1}^{n-1}\sum_{l=1}^{n-1}\mathbb{E}[r_{n-1,j}r_{n,n}\overline{r_{n-1,l}r_{n,n}}]\mathbb{E}[s_{j,n-1}\overline{s_{l,n-1}}]$
$=\frac{1}{n-1}\sum_{j=1}^{n-1}\mathbb{E}[r_{n-1,j}r_{n,n}\overline{r_{n-1,j}r_{n,n}}].$
Using Proposition 2.10 (see also Example 2.11), we have
$\mathbb{E}[r_{n-1,j}r_{n,n}\overline{r_{n-1,j}r_{n,n}}]=\begin{cases}\frac{1}{n^{2}(n+1)}&\text{if $j<n-1$};\\ \frac{n^{2}-n+2}{n^{2}(n+1)}&\text{if $j=n-1$},\end{cases}$
which gives
$\mathbb{E}\left[u_{n-1,n-1}u_{n,n}\overline{u_{n-1,n-1}u_{n,n}}\right]$ $=\frac{1}{n-1}\left((n-2)\cdot\frac{1}{n^{2}(n+1)}+\frac{n^{2}-n+2}{n^{2}(n+1)}\right)$
$=\frac{1}{(n-1)(n+1)},$
as desired.
This approach works for any integral of a monomial in $u_{ij}$ and their conjugates that has a single matching, resulting in a Weingarten function, and this provides a method to recompute the Weingarten function. It remains to identify the underpinning algebraic and combinatorial phenomena that allow one to rederive systematically the known results on Weingarten functions, which we leave as an open problem. For a related result, we also refer to Proposition 3.11.
3. A new recursive formula for Weingarten functions
3.1. A recursive formula
In this section, we rely on the achievements of the previous paragraphs to uncover a new relation between the Weingarten functions $\operatorname{Wg}_{k,n}$ and $\operatorname{Wg}_{k,n-1}$. Let $\mathbb{C}[S_{k}]$ denote the group algebra of the symmetric group, which consists of all complex-valued functions on $S_{k}$. The group algebra is equipped with a convolution product defined as follows:
$(f_{1}*f_{2})(\pi)=\sum_{\sigma\in S_{k}}f_{1}(\sigma)f_{2}(\sigma^{-1}\pi)\qquad(\pi\in S_{k}).$
Here, $f_{1}$ and $f_{2}$ are elements of $\mathbb{C}[S_{k}]$. If $f_{1}$ and $f_{2}$ are class functions on $S_{k}$, which means that they are constant on each conjugacy class of $S_{k}$, then the convolution can also be written as
$(f_{1}*f_{2})(\pi)=\sum_{\sigma\in S_{k}}f_{1}(\sigma)f_{2}(\sigma\pi),$
and it follows that $f_{1}*f_{2}=f_{2}*f_{1}$. We use such substitutions without notice.
We introduce a new element of $\mathbb{C}[S_{k}]$, which has already appeared in Example 2.13.
Definition 3.1.
Let $k$ and $n$ be positive integers. We define the ascension function $\operatorname{as}_{k,n}$ in $\mathbb{C}[S_{k}]$ by
$\operatorname{as}_{k,n}(\sigma)=\sum_{t=0}^{\mathsf{f}(\sigma)}(-1)^{k-t}\binom{\mathsf{f}(\sigma)}{t}\frac{1}{n^{\uparrow(k-t)}}\qquad(\sigma\in S_{k}),$
where $\mathsf{f}(\sigma)$ is the number of fixed points in $\sigma$:
$\mathsf{f}(\sigma)=|\{i\in\{1,2,\dots,k\}\ |\ \sigma(i)=i\}|.$
Example 3.2.
For the identity permutation $e_{k}$ in $S_{k}$, we have
$\operatorname{as}_{k,n}(e_{k})=1-\frac{k}{n}+\frac{\binom{k}{2}}{n(n+1)}-\cdots+(-1)^{k}\frac{1}{n(n+1)\cdots(n+k-1)}.$
If $\sigma\in S_{k}$ has no fixed points, then
$\operatorname{as}_{k,n}(\sigma)=\frac{(-1)^{k}}{n(n+1)\cdots(n+k-1)}.$
Recall the Weingarten function already appeared in the introduction. Suppose that $k\leq n$. The Weingarten function for the unitary group $\mathrm{U}(n)$ is given by
$\operatorname{Wg}_{k,n}(\pi)=\mathbb{E}\left[u_{11}u_{22}\cdots u_{kk}\overline{u_{1\pi(1)}u_{2\pi(2)}\cdots u_{k\pi(k)}}\right],$
for each permutation $\pi\in S_{k}$, where $U=(u_{ij})$ is a Haar-distributed unitary matrix in $\mathrm{U}(n)$. The origin of the name of $\operatorname{as}_{k,n}$ is based on the following theorem.
Theorem 3.3.
Let $k$ and $n$ be positive integers and suppose $k+1\leq n$. Then we have the convolution identity
$\operatorname{Wg}_{k,n}=\operatorname{as}_{k,n}*\operatorname{Wg}_{k,n-1}.$
Remark 3.4.
Theorem 3.3 does not hold without the assumption $k+1\leq n$. In the case where $k+1>n$, the Weingarten function $\operatorname{Wg}_{k,n-1}$ on the right-hand side is not well defined. Although the Weingarten functions can still be extended in that case, the above formula does not hold. We will elaborate on these points in Section $3.$
The following lemma will be used for the proof of Theorem 3.3.
Lemma 3.5.
Let $X=(x_{1},\dots,x_{n})^{\mathtt{T}}$ be a uniform vector on the unit sphere of $\mathbb{C}^{n}$. Then, for $k\leq n$ and any $\sigma\in S_{k}$, we have
$\operatorname{as}_{k,n}(\sigma)=\mathbb{E}\left[\prod_{i=1}^{k}\left(\delta_{i,\sigma(i)}-x_{i}\overline{x_{\sigma(i)}}\right)\right].$
Proof.
We denote the set of all fixed points of $\sigma$ by
$\operatorname{Fix}(\sigma)=\{i\in\{1,2,\dots,k\}\ |\ \sigma(i)=i\}.$
Expanding the factors, we have
$\mathbb{E}\left[\prod_{i=1}^{k}\left(\delta_{i,\sigma(i)}-x_{i}\overline{x_{\sigma(i)}}\right)\right]$
$=(-1)^{k}\,\mathbb{E}\left[\prod_{i\in\operatorname{Fix}(\sigma)}(|x_{i}|^{2}-1)\cdot\prod_{j\in\operatorname{Fix}(\sigma)^{c}}x_{j}\overline{x_{\sigma(j)}}\right]$
$=\sum_{s=0}^{\mathsf{f}(\sigma)}(-1)^{k+\mathsf{f}(\sigma)-s}\sum_{\{i_{1}<\cdots<i_{s}\}\subset\operatorname{Fix}(\sigma)}\mathbb{E}\left[|x_{i_{1}}|^{2}\cdots|x_{i_{s}}|^{2}\prod_{j\in\operatorname{Fix}(\sigma)^{c}}|x_{j}|^{2}\right],$
where $\operatorname{Fix}(\sigma)^{c}:=\{1,2,\dots,k\}\setminus\operatorname{Fix}(\sigma)$ is invariant under $\sigma$. By Lemma 2.2, each factor in the last sum is equal to $1/n^{\uparrow(s+k-\mathsf{f}(\sigma))}$. Therefore, we have
$\mathbb{E}\left[\prod_{i=1}^{k}\left(\delta_{i,\sigma(i)}-x_{i}\overline{x_{\sigma(i)}}\right)\right]$ $=\sum_{s=0}^{\mathsf{f}(\sigma)}(-1)^{k+\mathsf{f}(\sigma)-s}\binom{\mathsf{f}(\sigma)}{s}\frac{1}{n^{\uparrow(s+k-\mathsf{f}(\sigma))}}$
$=\sum_{t=0}^{\mathsf{f}(\sigma)}(-1)^{k-t}\binom{\mathsf{f}(\sigma)}{t}\frac{1}{n^{\uparrow(k-t)}}.$
Proof of Theorem 3.3.
Let the matrices $U=(u_{ij})$, $R=(r_{ij})$, and $V=(v_{ij})$ be as stated in Lemma 2.1. Then, for $i,j<n$, the $(i,j)$-th element of $U$ is given by
$u_{ij}=\sum_{p=1}^{n-1}r_{ip}v_{pj}.$
Let $\pi\in S_{k}$, and we consider
$\operatorname{Wg}_{k,n}(\pi)=\mathbb{E}\left[u_{1,1}u_{2,2}\cdots u_{k,k}\overline{u_{\pi(1),1}u_{\pi(2),2}\cdots u_{\pi(k),k}}\right].$
Since we are currently assuming $k+1\leq n$, the elements of $U$ in the $n$-th row or $n$-th column do not appear in this equation. By the independence of $R$ and $V$, we have
$\operatorname{Wg}_{k,n}(\pi)$ $=\sum_{p_{1}=1}^{n-1}\cdots\sum_{p_{k}=1}^{n-1}\sum_{q_{1}=1}^{n-1}\cdots\sum_{q_{k}=1}^{n-1}\mathbb{E}\left[r_{1,p_{1}}\cdots r_{k,p_{k}}\overline{r_{\pi(1),q_{1}}\cdots r_{\pi(k),q_{k}}}\right]$
$\qquad\times\mathbb{E}\left[v_{p_{1},1}\cdots v_{p_{k},k}\overline{v_{q_{1},1}\cdots v_{q_{k},k}}\right].$
Next, we apply the Weingarten calculus (1.1) for $\mathrm{U}(n-1)$ to the last factor:
$\mathbb{E}\left[v_{p_{1},1}\cdots v_{p_{k},k}\overline{v_{q_{1},1}\cdots v_{q_{k},k}}\right]$
$=\sum_{\tau\in S_{k}}\delta(p_{\tau^{-1}(1)},q_{1})\cdots\delta(p_{\tau^{-1}(k)},q_{k})\operatorname{Wg}_{k,n-1}(\tau).$
Substituting this into the preceding equation, we obtain
$\operatorname{Wg}_{k,n}(\pi)$
$=\sum_{\tau\in S_{k}}\operatorname{Wg}_{k,n-1}(\tau)\sum_{p_{1}=1}^{n-1}\cdots\sum_{p_{k}=1}^{n-1}\mathbb{E}\Big[r_{1,p_{1}}\cdots r_{k,p_{k}}\overline{r_{\pi\tau(1),p_{1}}\cdots r_{\pi\tau(k),p_{k}}}\Big],$
or $\operatorname{Wg}_{k,n}(\pi)=(\widetilde{\operatorname{as}}*\operatorname{Wg}_{k,n-1})(\pi)$ with
$\widetilde{\operatorname{as}}(\sigma):=\sum_{p_{1}=1}^{n-1}\cdots\sum_{p_{k}=1}^{n-1}\mathbb{E}\Big[r_{1,p_{1}}\cdots r_{k,p_{k}}\overline{r_{\sigma(1),p_{1}}\cdots r_{\sigma(k),p_{k}}}\Big]\qquad(\sigma\in S_{k}).$
The final step in the proof is to prove that $\widetilde{\operatorname{as}}(\sigma)$ is equal to $\operatorname{as}_{k,n}(\sigma)$, as defined in Definition 3.1. We start with the expression for $\widetilde{\operatorname{as}}(\sigma)$:
$\widetilde{\operatorname{as}}(\sigma)=\mathbb{E}\left[\left(\sum_{p_{1}=1}^{n-1}r_{1,p_{1}}\overline{r_{\sigma(1),p_{1}}}\right)\cdots\left(\sum_{p_{k}=1}^{n-1}r_{k,p_{k}}\overline{r_{\sigma(k),p_{k}}}\right)\right].$
Since $R=(r_{ij})$ is an $n\times n$ unitary matrix, we can analyze the sums involved. Specifically, we have:
$\sum_{p=1}^{n-1}r_{i,p}\overline{r_{\sigma(i),p}}=\sum_{p=1}^{n}r_{i,p}\overline{r_{\sigma(i),p}}-r_{i,n}\overline{r_{\sigma(i),n}}=\delta_{i,\sigma(i)}-r_{i,n}\overline{r_{\sigma(i),n}}$
for all $1\leq i\leq k$. From (2.1), we know that $r_{i,n}=x_{i}$. Substituting this into the above equation, we obtain:
$\widetilde{\operatorname{as}}(\sigma)=\mathbb{E}\left[\prod_{i=1}^{k}\big(\delta_{i,\sigma(i)}-x_{i}\overline{x_{\sigma(i)}}\big)\right].$
This equation, combined with Lemma 3.5, shows that $\widetilde{\operatorname{as}}(\sigma)=\operatorname{as}_{k,n}(\sigma)$. This completes the proof of Theorem 3.3. ∎
3.2. Some properties for ascension functions
In this subsection, we investigate some properties of the function $\operatorname{as}_{k,n}:S_{k}\to\mathbb{C}$. To do this, we review some fundamental properties of the Weingarten function, refer to [CMN22] and its references for details.
Consider the function $\operatorname{G}_{k,n}:S_{k}\to\mathbb{C}$ defined by
$\operatorname{G}_{k,n}(\pi)=n^{\kappa(\pi)}\qquad(\pi\in S_{k}),$
where $\kappa(\pi)$ is the number of cycles in the cycle decomposition of $\pi$. This function is invertible if and only if $k\leq n$, and in this case, the inverse function of $\operatorname{G}_{k,n}$ is the Weingarten function $\operatorname{Wg}_{k,n}$:
$\operatorname{G}_{k,n}*\operatorname{Wg}_{k,n}=\operatorname{Wg}_{k,n}*\operatorname{G}_{k,n}=\delta_{e_{k}}.$
Here $\delta_{e_{k}}:S_{k}\to\{0,1\}$ is the Dirac function at the identity permutation $e_{k}$:
$\delta_{e_{k}}(\sigma)=\begin{cases}1&\text{if $\sigma=e_{k}$,}\\ 0&\text{otherwise}.\end{cases}$
From Theorem 3.3, it is immediate to see the formula
(3.1) $\operatorname{as}_{k,n}=\operatorname{G}_{k,n-1}*\operatorname{Wg}_{k,n}=\operatorname{Wg}_{k,n}*\operatorname{G}_{k,n-1}$
if $k+1\leq n$.
Next, we consider character expansions for these functions. A partition of a positive integer $k$ is a weakly decreasing sequence $\lambda=(\lambda_{1},\lambda_{2},\dots,\lambda_{l})$ of positive integers such that the sum of all parts $\lambda_{i}$ is equal to $k$. In this case, we write $\lambda\vdash k$ and $\ell(\lambda)=l$. For a partition $\lambda\vdash k$ and a complex number $z$, we define
$(z\uparrow\lambda)=\prod_{i=1}^{\ell(\lambda)}\prod_{j=1}^{\lambda_{i}}(z+j-i).$
Denote by $\chi^{\lambda}$ the irreducible character of $S_{k}$ corresponding to $\lambda$. Moreover, we define $f^{\lambda}:=\chi^{\lambda}(e_{k})$, the value of $\chi^{\lambda}$ at the identity permutation. Using these notations, we have the following expansions:
(3.2) $\operatorname{G}_{k,n}=\frac{1}{k!}\sum_{\lambda\vdash k}f^{\lambda}(n\uparrow\lambda)\chi^{\lambda},\qquad\operatorname{Wg}_{k,n}=\frac{1}{k!}\sum_{\lambda\vdash k}\frac{f^{\lambda}}{(n\uparrow\lambda)}\chi^{\lambda}.$
Note that we need to assume $k\leq n$ in the latter equation to ensure that the factor $(n\uparrow\lambda)$ of all terms does not equal 0. From these formulas, we can easily obtain the character expansion of the ascension function using the orthogonality of irreducible characters
$\chi^{\lambda}*\chi^{\mu}=\delta_{\lambda,\mu}\frac{k!}{f^{\lambda}}\chi^{\lambda}\qquad(\lambda,\mu\vdash k).$
Indeed, (3.1) implies that
(3.3) $\operatorname{as}_{k,n}=\frac{1}{k!}\sum_{\lambda\vdash k}f^{\lambda}\left(\prod_{i=1}^{d(\lambda)}\frac{n-1-(\lambda_{i}^{\prime}-i)}{n-(\lambda_{i}-i)}\right)\chi^{\lambda},$
where $(\lambda^{\prime}_{1},\lambda_{2}^{\prime},\dots)$ represents the conjugate partition of $\lambda$, obtained by transposing the partition when viewed as a Young diagram along its diagonal. Additionally, $d(\lambda)=|\{i\geq 1\ |\ \lambda_{i}\geq i\}|$ denotes the length of the diagonal in the Young diagram $\lambda$. Here, the ratio in the coefficient can be derived from the following identity:
$\frac{((n-1)\uparrow\lambda)}{(n\uparrow\lambda)}=\prod_{i=1}^{d(\lambda)}\frac{n-1-(\lambda_{i}^{\prime}-i)}{n-(\lambda_{i}-i)},$
which can be obtained through simple factor cancellation.
Finally, we discuss the inverse of the ascension function.
Definition 3.6.
Let $k$ and $n$ be positive integers. We assume $k\leq n$ and define the descension function $\operatorname{ds}_{k,n}$ in $\mathbb{C}[S_{k}]$ by
$\operatorname{ds}_{k,n}(\sigma)=(\operatorname{sgn}\sigma)\sum_{t=0}^{\mathsf{f}(\sigma)}\binom{\mathsf{f}(\sigma)}{t}\frac{1}{n^{\downarrow(k-t)}}\qquad(\sigma\in S_{k}),$
where $n^{\downarrow t}$ is the falling factorial defined by
$n^{\downarrow t}=\begin{cases}n(n-1)\cdots(n-t+1)&\text{for $t=1,2,\dots$;}\\ 1&\text{for $t=0$}.\end{cases}$
The falling factorial term imposes the assumption $k\leq n$.
Proposition 3.7.
Let $k$ and $n$ be positive integers, and we assume $k\leq n$. Then it holds that
(3.4) $\operatorname{as}_{k,n+1}*\operatorname{ds}_{k,n}=\delta_{e_{k}}.$
Furthermore, it holds that
$\operatorname{Wg}_{k,n}=\operatorname{ds}_{k,n}*\operatorname{Wg}_{k,n+1}.$
Proof.
The second formula follows from Theorem 3.3 and (3.4). To show the first formula (3.4), we extend the range of the parameter of the functions $\operatorname{G}_{k,n}$, $\operatorname{Wg}_{k,n}$, and $\operatorname{as}_{k,n}$ from a positive integer $n$ (where $k\leq n$) to a complex number $z$, as referenced in [CŚ06]. We define $\operatorname{G}_{k,z}$ by
$\operatorname{G}_{k,z}(\sigma):=z^{\kappa(\sigma)}\quad(\sigma\in S_{k}).$
This function has a trivial symmetry:
$\operatorname{G}_{k,-z}(\sigma)=(-1)^{\kappa(\sigma)}\operatorname{G}_{k,z}(\sigma)=(-1)^{k}\operatorname{sgn}(\sigma)\operatorname{G}_{k,z}(\sigma)$
since $\operatorname{sgn}\sigma=(-1)^{k-\kappa(\sigma)}$. It is known that $\operatorname{G}_{k,z}$ is invertible in the group algebra $\mathbb{C}[S_{k}]$ if and only if $z\not\in\{0,1,\dots,k-1\}$. This fact can be confirmed from (3.2). In this case, we denote by $\operatorname{Wg}_{k,z}$ the inverse of $\operatorname{G}_{k,z}$. It also satisfies the symmetry relation $\operatorname{Wg}_{k,-z}(\sigma)=(-1)^{k}\operatorname{sgn}(\sigma)\operatorname{Wg}_{k,z}(\sigma)$.
Similarly, we define
$\operatorname{as}_{k,z}(\sigma):=(\operatorname{G}_{k,z-1}*\operatorname{Wg}_{k,z})(\sigma)=\sum_{t=0}^{\mathsf{f}(\sigma)}(-1)^{k-t}\binom{\mathsf{f}(\sigma)}{t}\frac{1}{z^{\uparrow(k-t)}},$
based on (3.1). This function is invertible if and only if $z\not\in\{1,\dots,k\}$. In particular, if $k\leq n$ then $\operatorname{as}_{k,n+1}$ is invertible. Since $\operatorname{G}_{k,z}$ and $\operatorname{Wg}_{k,z}$ are inverses of each other for all $z\not\in\{0,1,\dots,k-1\}$, when we denote the inverse of $\operatorname{as}_{k,z+1}$ as $\operatorname{as}^{-1}_{k,z+1}$, we can transform it as follows using symmetries:
$\operatorname{as}^{-1}_{k,z+1}(\sigma)$ $=(\operatorname{Wg}_{k,z}*\operatorname{G}_{k,z+1})(\sigma)$
$=(\operatorname{sgn}\sigma)(\operatorname{Wg}_{k,-z}*\operatorname{G}_{k,-z-1})(\sigma)$
$=(\operatorname{sgn}\sigma)\operatorname{as}_{k,-z}(\sigma).$
Therefore, the inverse of $\operatorname{as}_{k,n+1}$ is $\operatorname{sgn}*\operatorname{as}_{k,-n}$, which is equal to $\operatorname{ds}_{k,n}$. Indeed, we have:
$\operatorname{sgn}(\sigma)\operatorname{as}_{k,-n}(\sigma)$ $=\operatorname{sgn}(\sigma)\sum_{t=0}^{\mathsf{f}(\sigma)}(-1)^{k-t}\binom{\mathsf{f}(\sigma)}{t}\frac{1}{(-n)^{\uparrow(k-t)}}$
$=\operatorname{sgn}(\sigma)\sum_{t=0}^{\mathsf{f}(\sigma)}\binom{\mathsf{f}(\sigma)}{t}\frac{1}{n^{\downarrow(k-t)}}$
$=\operatorname{ds}_{k,n}(\sigma).$
Remark 3.8.
We have proved Theorem 3.3 using random matrices; however, since it is an algebraic equation, it would be interesting to look for an algebraic proof.
3.3. Weingarten functions in the case of low dimensions
In the previous discussions, we often assumed that $k\leq n$. This is due to the fact that the function $\operatorname{G}_{k,n}(\sigma)=n^{\kappa(\sigma)}$ on $S_{k}$ is invertible only when $k\leq n$, and the inverse element is $\operatorname{Wg}_{k,n}$. The Weingarten calculus itself (1.1) holds even when $k>n$; however, in that case, it is necessary to replace $\operatorname{Wg}_{k,n}$ with the next $w_{k,n}$. The function $w_{k,n}$ on $S_{k}$ is a class function satisfying the following equation:
(3.5) $\operatorname{G}_{k,n}*w_{k,n}*\operatorname{G}_{k,n}=\operatorname{G}_{k,n}.$
Such functions $w_{k,n}$ are not uniquely defined if $k>n$, but they are also referred to as Weingarten functions. When $k\leq n$, $w_{k,n}$ is uniquely determined and coincides with $\operatorname{Wg}_{k,n}$.
Remark 3.9.
Assume that $k>n$. If we also assume $w_{k,n}*\operatorname{G}_{k,n}*w_{k,n}=w_{k,n}$ in addition to (3.5), then $w_{k,n}$ is uniquely determined. (This condition was insufficient in [Mat13].) If we denote such a $w_{k,n}$ specifically as $W_{k,n}$, then $W_{k,n}$ is concretely given by
$W_{k,n}=\frac{1}{k!}\sum_{\begin{subarray}{c}\lambda\vdash k\\ \ell(\lambda)\leq n\end{subarray}}\frac{f^{\lambda}}{(n\uparrow\lambda)}\chi^{\lambda},$
which includes a slight restriction $\ell(\lambda)\leq n$ added to (3.2). Here, $\ell(\lambda)=|\{i\geq 1\ |\ \lambda_{i}>0\}|$ is the length of a partition $\lambda=(\lambda_{1},\lambda_{2},\dots)$. As already mentioned, the Weingarten formula (1.1) holds for any $w_{k,n}$ that satisfies (3.5), not just for this special $W_{k,n}$ – we refer to [CFM25] for further details.
Unfortunately, Theorem 3.3 does not hold for functions $w_{k,n}$ when $k\geq n$. Indeed, let $k=n=2$ and $\sigma=e_{2}\in S_{2}$. If we apply Theorem 3.3 formally in this case, it is expected that the following holds:
(3.6) $w_{2,2}(e_{2})\stackrel{{\scriptstyle?}}{{=}}(\operatorname{as}_{2,2}*w_{2,1})(e_{2})=\sum_{\pi\in S_{2}}\operatorname{as}_{2,2}(\pi)w_{2,1}(\pi).$
It is easy to see that $w_{2,2}(e_{2})=\operatorname{Wg}_{2,2}(e_{2})=\frac{1}{3}$, while Definition 3.1 and (3.5) imply $\sum_{\pi\in S_{2}}\operatorname{as}_{2,2}(\pi)w_{2,1}(\pi)=\frac{1}{6}\sum_{\pi\in S_{2}}w_{2,1}(\pi)=\frac{1}{12}.$ Therefore, for the case when $k\geq n$, a different equation corresponding to Theorem 3.3 needs to be sought. In the general case, it is not possible, but for Theorem 3.3 when $k=n$, we could obtain a similar expression as follows.
Theorem 3.10.
For $\pi\in S_{n}$, we have
$\operatorname{Wg}_{n,n}(\pi)=\sum_{\tau\in S_{n-1}}a_{n}(\pi,\tau)\operatorname{Wg}_{n-1,n-1}(\tau),$
where the function $a_{n}:S_{n}\times S_{n-1}\to\mathbb{C}$ is defined by
(3.7) $a_{n}(\pi,\tau)=\sum_{t=0}^{\mathsf{f}(\pi,\tau)}(-1)^{n-t+1}\binom{\mathsf{f}(\pi,\tau)}{t}\frac{1}{n^{\uparrow(n-t)}}$
and $\mathsf{f}(\pi,\tau)=|\{i\in\{1,2,\dots,n-1\}\ |\ \pi(\tau(i))=i\}|$.
In this theorem, it should be noted that $a_{n}(\pi,\tau)$ is slightly different from $\operatorname{as}_{n,n}(\pi\tau)$.
According to this theorem, the incorrect formula (3.6) can be corrected as follows:
$\operatorname{Wg}_{2,2}(e_{2})=a_{2}(e_{2},e_{1})\operatorname{Wg}_{1,1}(e_{1}),$
both of which are equal to $\frac{1}{3}$. As another example of the theorem, the following equation holds true:
$\operatorname{Wg}_{3,3}(e_{3})=a_{3}(e_{3},e_{2})\operatorname{Wg}_{2,2}(e_{2})+a_{3}(e_{3},(1\,2))\operatorname{Wg}_{2,2}((1\,2)).$
When substituting specific values, this equation becomes:
$\frac{7}{120}=\frac{11}{60}\cdot\frac{1}{3}+\frac{1}{60}\cdot\left(-\frac{1}{6}\right).$
Proof of Theorem 3.10.
The proof is almost the same as that of Theorem 3.3. we will focus on explaining the different parts. Let the matrices $U=(u_{ij})$, $R=(r_{ij})$, and $V=(v_{ij})$ be as stated in Lemma  2.1. Let $\pi\in S_{n}$, and we consider
$\operatorname{Wg}_{n,n}(\pi)=\mathbb{E}\left[u_{1,1}\cdots u_{n-1,n-1}u_{n,n}\overline{u_{\pi(1),1}\cdots u_{\pi(n-1),n-1}u_{\pi(n),n}}\right].$
We substitute
$u_{ij}=\begin{cases}\sum_{p=1}^{n-1}r_{i,p}v_{p,j}&\text{if $1\leq i\leq n$ and $1\leq j\leq n-1$};\\ r_{i,n}&\text{if $1\leq i\leq n$ and $j=n$}.\end{cases}$
After some direct calculations, we have:
$\operatorname{Wg}_{n,n}(\pi)=\sum_{\tau\in S_{n-1}}\tilde{a}(\pi,\tau)\operatorname{Wg}_{n-1,n-1}(\tau),$
where
$\tilde{a}(\pi,\tau)=\mathbb{E}\left[r_{1,p_{1}}\cdots r_{n-1,p_{n-1}}r_{n,n}\overline{r_{\pi(\tau(1)),p_{1}}\cdots r_{\pi(\tau(n-1)),p_{n-1}}r_{\pi(n),n}}\right].$
As with the proof of Theorem 3.3, we obtain the following expression:
$\tilde{a}(\pi,\tau)=\mathbb{E}\left[\prod_{i=1}^{n-1}\left(\delta(i,\pi(\tau(i)))-x_{i}\overline{x_{\pi(\tau(i))}}\right)\times x_{n}\overline{x_{\pi(n)}}\right].$
From here, we mimic the proof of Lemma 3.5. Define
$F=\{i\in\{1,2,\dots,n-1\}\ |\ \pi(\tau(i))=i\}|,$
and put $F^{\prime}=\{1,2,\dots,n\}\setminus F$. Note that $F^{\prime}$ is not the complement in $\{1,2,\dots,n-1\}$. Then, we have:
$\tilde{a}(\pi,\tau)$ $=(-1)^{n-1}\mathbb{E}\left[\prod_{i\in F}(|x_{i}|^{2}-1)\cdot\prod_{j\in F^{\prime}}|x_{j}|^{2}\right]$
$=\sum_{s=0}^{|F|}(-1)^{n-1+|F|-s}\sum_{\{i_{1}<\cdots<i_{s}\}\subset F}\mathbb{E}\left[|x_{i_{1}}|^{2}\cdots|x_{i_{s}}|^{2}\prod_{j\in F^{\prime}}|x_{j}|^{2}\right]$
$=\sum_{s=0}^{|F|}(-1)^{n-1+|F|-s}\binom{|F|}{s}\frac{1}{n^{\uparrow(s+n-|F|)}}$
$=\sum_{t=0}^{|F|}(-1)^{n-1+t}\binom{|F|}{t}\frac{1}{n^{\uparrow(n-t)}}=a_{n}(\pi,\tau),$
as desired. We used Lemma 2.2 in the third equality. ∎
Finally, let us assume $k<n$ and let $\pi\in S_{k}$. By repeatedly applying Theorem 3.3, we have
$\operatorname{Wg}_{k,n}(\pi)$ $=\sum_{\sigma_{n-1}\in S_{k}}\sum_{\sigma_{n-2}\in S_{k}}\cdots\sum_{\sigma_{k}\in S_{k}}\operatorname{as}_{k,n}(\pi\sigma_{n-1}^{-1})\,\operatorname{as}_{k,n-1}(\sigma_{n-1}\sigma_{n-2}^{-1})\,\cdots\,$
$\qquad\times\operatorname{as}_{k,k+1}(\sigma_{k+1}\sigma_{k}^{-1})\operatorname{Wg}_{k,k}(\sigma_{k}).$
Furthermore, Theorem 3.10 implies that, for each $\sigma\in S_{k}$
$\operatorname{Wg}_{k,k}(\sigma)=\sum_{\tau_{k-1}\in S_{k-1}}\sum_{\tau_{k-2}\in S_{k-2}}\cdots\sum_{\tau_{2}\in S_{2}}a_{k}(\sigma,\tau_{k-1})\,a_{k-1}(\tau_{k-1},\tau_{k-2})\,\cdots\,a_{2}(\tau_{2},e_{1}).$
By combining these, the Weingarten function $\operatorname{Wg}_{k,n}\in\mathbb{C}[S_{k}]$ with $k\leq n$ can be expressed using two types of functions: $\operatorname{as}_{k,*}(\sigma)$ and $a_{*}(\sigma,\tau)$.
Proposition 3.11.
Let $k$ and $n$ be positive integers, and suppose $k\leq n$. Then, for any $\pi\in S_{k}$, we have
$\operatorname{Wg}_{k,n}(\pi)$ $=\sum\operatorname{as}_{k,n}(\pi\sigma_{n-1}^{-1})\,\operatorname{as}_{k,n-1}(\sigma_{n-1}\sigma_{n-2}^{-1})\,\cdots\,\operatorname{as}_{k,k+1}(\sigma_{k+1}\sigma_{k}^{-1})$
$\qquad\times a_{k}(\sigma_{k},\tau_{k-1})\,a_{k-1}(\tau_{k-1},\tau_{k-2})\,\cdots\,a_{2}(\tau_{2},e_{1})$
summed over all sequences of permutations
$(\underbrace{\sigma_{n-1},\sigma_{n-2},\dots,\sigma_{k}}_{n-k},\tau_{k-1},\tau_{k-2},\dots,\tau_{2})\in\underbrace{S_{k}\times\cdots\times S_{k}}_{n-k}\times S_{k-1}\times S_{k-2}\times\cdots\times S_{2}.$
Here, $\operatorname{as}_{k,n}(\sigma)$ is defined in Definition 3.1, and $a_{n}(\sigma,\tau)$ is defined in (3.7).
This theorem allows the Weingarten function to be calculated sequentially starting from $n=1$.